Hi, I need answer for this question below.BR//Hassan suppose that net exports, NXt , are determined...

Hi,

I need answer for this question below.BR//Hassan

suppose that net exports, NXt , are determined by the following expression:

NXt Y¯ t = αNX − βNX(rt − r ∗ t ),

where Y¯ t is potential GDP and rt − r ∗ t is the difference between the domestic and the foreign real interest rate.

a. Explain the intution behind expression .

b. Modify the IS curve .

c. Explain how an increase in the domestic real interest rate affects short-run output in this economy.
Answer #1

a.NXt Y¯ t = αNX − βNX(rt − r ∗ t ). Where net exports = NXt, Y¯ t = potential GDP , rt is the domestic real interest rate and r ∗ t is the foreign real interest rate. When prices are sticky, there will be an increase in the nominal interest rate differential and hence nominal appreciation will be triggered by this increased rate. Consider an increase in the real interest rate (rt −r∗t) and so an increase in rt −r∗t triggers a real appreciation which leads to a decrease in net exports (NXt).

b. The IS curve is written as follows-

Y˜t = αC +αG +αI +αNX −1− βI(rt −r¯)−βNX (rt −r¯) +βNX (r∗t −r¯)

Where α = αC +αG +αI +αNX −1 +βNX (r∗t −r¯),

β = βI +βNX

The IS curve could be rewritten as follows -
Y˜t = α −β(rt −r¯)

where α ≡ αC +αG +αI +αEX −αIM −1 , ( C and t are constant fractions of potential output, I stands for Investment, EX and IM stands for exports and imports respectively).

c. In the short run, the increase in the domestic real interest rate decreases the domestic output and appreciates the current real exchange rate and the net exports of the domestic country increases only if the effect of lower output is stronger than the effect of increase in the real exchange rate.